Green Tractor needs to raise $30 to expand their distribution network. They will selling equity for cash to raise this money. If the underwriting spread they will pay is 8.8%, and they will have to pay $2 million in filing fees, how many dollars of equity do they need to sell?
Answer #1

Suppose Gross Proceed from sale of equity is $X. Net Proceed to company can be computed with following equation:

Net Proceed = Gross Proceed - Gross Proceed*Underwriting spread - Filing fees

30,000,000 = X - X*0.088 - 2,000,000

X-X*0.088 = 30,000,000+2,000,000

X*0.912 = 32,000,000

X = 32,000,000 0.912

X = $35.087.719.30

Thus, Green Tractor need to sell $35,087,719.30 of equity to get $30,000,000.

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.